OpenStudy (anonymous):

The edge of a cube was found to be 30 cm with a possible error in measurement of .1 cm. Use differentials to estimate the percentage error (to the nearest hundredth) in computing (a) the volume of the cube and (b) the surface area of the cube.

OpenStudy (anonymous):

V = a³ -> V' = 3 a² a' = 3*30²* .1 % error = V' * 100/ V = .. Similarly for Surface area: @alexkolyszko You should be fine from here :)